OpenStudy (anonymous):

Find the midpoint of the line segment joining the points p1 and p2 p1(4,2) p2(-5,4)

OpenStudy (anonymous):

midpoint of x = (x1 + x2)/2 = (4 + -5)/2 midpoint of y = (y1 + y2)/2 = (2 +4)/2

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you're welcome :)